A salesperson in an expensive clothing store earns $2600 per month base pay plus a 6.5%
commission on sales. One month she earns $5518.50. What were her sales for that month?

Answer by ikleyn(52781)

.

2600 + 0.065*x = 5518.50  dollars


Simplify and find x

       0.065x = 5518.50 - 2600

       0.065x =    2918.5

            x =    2918.5/0.065 = 44900 dollars.      ANSWER

Solved.

The equations are self-explanatory.